OpenStudy (cheesecakekitten):

WILL MEDAL AND FAN!!!! What is the ratio of a to b? 8 / 15 = 4a / (a + b) A. 11/ 15 B. 4 / 15 C. 2 / 13 D. 59 / 8

OpenStudy (cheesecakekitten):

Thanks in advance.

OpenStudy (gabebae):

Do you need the answer or do you want me to explain?

OpenStudy (cheesecakekitten):

I need to answer, not to be a bother..

OpenStudy (cheesecakekitten):

*the not to

OpenStudy (gabebae):

C

OpenStudy (gabebae):

You need to find a and b and put them as a/b

OpenStudy (cheesecakekitten):

Back, my browser crashed. Did I miss anything important?

OpenStudy (gabebae):

Nope welcome back

OpenStudy (cheesecakekitten):

Here's another one: One a soccer team, 11 out of 17 players surveyed say they had two or more siblings. The league has 850 players. Which proportion could be used to predict the number of players in the league that have two or more siblings? A. x / 11 = 17 / 850 B. x / 17 = 11 / 850 C. x / 850 = 17 / 11 D. x / 850 = 11 / 17

OpenStudy (cheesecakekitten):

Take your time. :)

OpenStudy (cheesecakekitten):

@mathmath333 @Mehek14 @TheSmartOne help please?

OpenStudy (gabebae):

Oh i'm sorry I didn't see it

OpenStudy (cheesecakekitten):

Oh that's fine. :)

OpenStudy (gabebae):

It would be D. You would line up the ratio for the soccer team and since the league has 850 players that would be the number on the bottom 

OpenStudy (gabebae):

Total number of soccer players is 17 and total numbers of players in the league is 850 so those would line up

OpenStudy (cheesecakekitten):

Thanks a bunch, I have like two more but you don't need to help on them if you don't want too. :)

OpenStudy (gabebae):

sure